OpenStudy (anonymous):

Hey guys, just a simple question. I know that you CAN'T take the square root of a negative number, but is this possible, taking the square root of -x?

hartnn (hartnn):

do you know about imaginary number \(i=\sqrt{-1}\) ?

OpenStudy (anonymous):

umm, i cant write the equation bc the equation button is not working for me. I have heard of them, but never used them. I know they have to do with a negative under the square root.

OpenStudy (anonymous):

but im working on a question and i got -x under a square root, is it possible?

hartnn (hartnn):

yes, ofcourse it is possible but if imaginar or complex numbers are not yet introduced to you, you should not get an example of imaginary number. can i have the original question, just to make sure ?

hartnn (hartnn):

bdw, \(\sqrt{-x} = \sqrt{-1}\sqrt{x}=i\sqrt{x}\) if x is positive and \(\sqrt{-x} = \sqrt{|x|}\) if x is negative

OpenStudy (anonymous):

umm the equation button is not working for me, i think i will just ask my prof